# DKDS: A Benchmark Dataset of Degraded Kuzushiji Documents with Seals for Detection and Binarization

### v1

@@ 5.3.2 Visualization
We present two groups of result images generated by different methods, along with their corresponding input and ground-truth images, as shown in Figs. 11 and 12. The input images contain various forms of degradation, including red seals, paper damage, stains, and ink fading, while the ground-truth images remove all such interference, retaining only the black Kuzushiji characters.
Among the traditional methods, Niblack’s algorithm [niblack1985introduction] produced the poorest results, with much noise remaining in the background. Otsu’s method [otsu1979threshold] achieved better separation between text and background but still retained some noise. Sauvola’s method [sauvola2000adaptive] performed relatively better but failed to completely remove seal interference.
When combined with K-means clustering [mcqueen1967some] preprocessing, all three traditional methods achieved improved performance, further reducing seal marks, though minor residues remained.
In contrast, the GAN-based method [goodfellow2020generative] obtained the best results, effectively removing most noise and seal interference while maintaining excellent text legibility.
@@ 6 Conclusion
Binarizing degraded Kuzushiji document images presents significant challenges, particularly when seals overlap with the characters. Effectively removing seal interference while preserving or restoring character details remains a critical problem. Notably, no publicly available dataset currently contains Kuzushiji documents with seals, limiting the development of methods for this task. To address this gap, we constructed the Degraded Kuzushiji Documents with Seals (DKDS) dataset by collecting and integrating pre-modern Kuzushiji document images with high-resolution images of ancient seals. We detail the construction process of the DKDS dataset and define two task tracks based on it: (1) text and seal detection, and (2) document binarization. Furthermore, we provide baseline evaluation results for several classical methods to facilitate future research.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Numbers 25H01242.
@@
Rui-Yang Ju: Conceptualization, Data Curation, Formal Analysis, Methodology, Writing – Original Draft Preparation, Writing – Review & Editing; Kohei Yamashita: Data Curation, Investigation, Writing – Review & Editing; Hirotaka Kameko: Project Administration, Resources, Writing – Review & Editing; Shinsuke Mori: Funding Acquisition, Resources, Supervision, Writing – Review & Editing.

### v2

@@ 5.3.2 Visualization
We present visualized results generated by different methods, along with their corresponding input and ground-truth images, as shown in Fig. 11. The input images contain various forms of degradation, including red seals, paper damage, stains, and ink fading, while the ground-truth images remove all such interference, retaining only the black Kuzushiji characters.
Among the traditional methods, Niblack’s algorithm [niblack1985introduction] produces the poorest results, with much noise remaining in the background. Otsu’s method [otsu1979threshold] achieves better separation between text and background but still retains some noise. Sauvola’s method [sauvola2000adaptive] performs relatively better but failed to completely remove seal interference.
Furthermore, as shown in Table 5, the incorporation of K-means clustering does not lead to a substantial performance improvement when documents contain a large number of seals.
In contrast, deep learning-based methods, including those proposed by Suh et al. [suh2022two], Ju et al. [ju2024three], and our method, achieve markedly better results by effectively suppressing noise and seal interference while preserving high text legibility.
5.4 OCR Evaluation
To quantitatively assess OCR performance on degraded and binarized documents, we employ the “miwo” app to recognize Kuzushiji characters. OCR is conducted on combined sets consisting of the training set and each version of the test set, with each evaluation involving a total of 50 document images.
The experimental results show that binarization leads to a substantial reduction in character error rate (CER). On the combined set of 50 images from the training set and the Testing-E set, the CER decreases by 27.6% in relative terms after binarization. Similarly, on the combined set of 50 images from the training set and the Testing-D set, the CER is reduced by 34.9%. Specifically, for the training set and Testing-E set, the CER of the binarized documents is 1.497%, compared with 2.067% for the original degraded documents. For the training set and Testing-D set, the CER of the original degraded documents reaches 2.301%, which is also markedly reduced after binarization (CER = 1.497%).
Notably, both the raw Kuzushiji document images [genjimonogatari] and the “miwo” OCR app [clanuwat2021miwo] are provided by the Center for Open Data in the Humanities (CODH), which may partially account for the strong recognition performance observed on these documents.
@@ 6 Conclusion
Binarizing degraded Kuzushiji document images presents significant challenges, particularly when seals overlap with the characters. Effectively removing seal interference while preserving or restoring character details remains a critical problem. Notably, no publicly available dataset currently contains Kuzushiji documents with seals, limiting the development of methods for this task. To address this gap, we constructed the Degraded Kuzushiji Documents with Seals (DKDS) dataset by collecting and integrating pre-modern Kuzushiji document images with high-resolution images of ancient seals. We detail the construction process of the DKDS dataset and define two task tracks based on it: (1) text and seal detection, and (2) document binarization. Furthermore, we provide baseline evaluation results for several classical and deep learning-based methods to facilitate future research.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Number 25H01242, and JST SPRING Grant Number JPMJSP2110.
@@
Rui-Yang Ju: Conceptualization, Data Curation, Formal Analysis, Methodology, Writing – Original Draft Preparation, Writing – Review & Editing; Kohei Yamashita: Methodology, Data Curation, Investigation, Writing – Review & Editing; Hirotaka Kameko: Project Administration, Resources, Writing – Review & Editing; Shinsuke Mori: Funding Acquisition, Resources, Supervision, Writing – Review & Editing.